40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1471     Entry time: Thu Aug 14 15:23:36 2014
Author: Emily, Evan 
Type: Notes 
Category: optic 
Subject: AOM fiber noise cancellation  

 New setup for fiber phase noise cancellation with one AOM

 
We re-did mode-matching calculations and replaced the lenses before the fiber input in order to optimize the amount of power that comes out of the fiber.  The waist coming out of the PMC is 370 microns.  Following the PMC are the following lenses: placed 7 inches away is a PLCX-25.4.128.8-UV-1064 with a focal length of 250mm, placed 29 inches away is a PLCX-25.4-64.4-C-1064 with a focal length of 125 mm, and placed 35 inches away is a KBX052 with a focal length of 50.2mm. This yields a waist of 69 microns going into the fiber.  Going into the fiber is about 1.1 mw and coming out is approximately 500 micro watts.  We replaced the VCO driver since it was not driving the AOM and had a deformed signal.  Now we are using a Marconi and low-noise amplifier to drive the AOM.  We also replaced the AOM with an Isomet AOM 1205c-843.  
 
We re-did mode-matching calculations into the AOM and to the mirror.  After the fiber output is a waist of 50 microns.  Placed 2 inches away is a: PLCX-25.4-33.7-UV-1064 with a focal length of 50mm, placed 10 inches away is a: PLCX-25.4-77.3-UV-1064 with a focal length of 150mm and placed 18 inches away is a: PLCX-25.4-36.1-UV-1064 with a focal length of 70mm.  The first two lenses before the AOM yield a was it of 150 microns going through the AOM (recommended waist from the Isomet AOM 1205c-843 manual) and the third lens yields a waist of 156 microns at the mirror.  We used a beam dump to block the zeroth order beam, so the only the first order beam is double passed through the fiber.  
 
We are using the same setup to beat the double passed beam with the original beam onto a new focus 1811 photodiode.  The original beam has a power of 850 micro-watts and the double-passed beam has a power of 10 micro-watts. While the efficiency can be improved, for now we will work with what we have in order to prove that our new setup with 1 AOM will cancel the noise in the system.  
 Final_AOM_Setup.pdf
In this setup, we lock the optical beat to the marconi in a PLL.  
The AC signal optical beat fluctuation was 198-428mV.
Once the optical beat was locked to the marconi, we measured the error signal and control signal.  We also measured the control signal without cancellation to make sure that this works.  In order to do the measurement without cancellation, we locked the marconi to the optical beat.  We also measured the open loop transfer function with and without cancellation.  The following data was obtained: 
 
 Unknown-4
Unknown-5
 
 
 
ELOG V3.1.3-